A board 1.2 m long is cut into two so that the length of the longer piece is 15 cm long than twice the length of the shorter piece. How long is the shorter piece?
Answer by mananth(16946)

let shorter piece be x
longer one will be 2x+15
..
2x+15 +x=120 cm
3x+15=120
-15 -15
3x=105
/3
x= 35 cm
the other will be 2x+15
= 2*35+15
=85 cm
..
m.ananth@hotmail.ca